OpenStudy (tahjna):

Find the slope m of the tangent line to the parabola f(x)=x2+2x at the point (4, 24) by using the definition: m=limh→0f(a+h)−f(a)/(h) m= Find the equation for the tangent line at the point (4, 24), and write it in the form y=mx+b y = m x + b where m = and b =

OpenStudy (mathmale):

First: some housekeeping: Please express exponentiation using the symbol " ^ ." Unclear: f(x)=x2+2x Clear: f(x)=x^2+2x

OpenStudy (mathmale):

Definition of derivative: f'(x)= (limit as h approaches zero) of \[\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (mathmale):

What is f(x+h)? Write out all terms. What is f(x+h)-f(x)? Simplifly your result. What is \[\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (mathmale):

After all possible simplifications, let h approach zero. Be sure to label your result with f '(x).